OpenStudy (hellomynameiskam):

Give the value of x that makes this equation true. (3)(–4)(x) = 0 A. 1 B. –3 C. 0 D. 4 i think its 0. am I correct?

OpenStudy (tkhunny):

Use the multiplication rule. What is the answer? In order for a product to be zero (0), one of the factors must be zero. Can 3 ever be zero? Can -4 ever be zero? Can x ever be zero?

OpenStudy (gabebae):

The answer is zero

OpenStudy (hellomynameiskam):

They could be zero's cause 3*0=0 -4*0=0 and x can be a 0

OpenStudy (alexandervonhumboldt2):

correct

OpenStudy (alexandervonhumboldt2):

0/a=0 always

OpenStudy (alexandervonhumboldt2):

thus 0/-12=0 as well

OpenStudy (hellomynameiskam):

i have like a few more questions left, can u guys help me?

OpenStudy (gabebae):

Hah I wasn't giving out the answer I was saying the (3)(–4)(x) EQUALS ZERO they're trying to solve for x